Suppose there are two independent economic factors, M1 and M2. The risk-free rate is 5%, and all stocks have independent firm-specific components with a standard deviation of 44%. Portfolios A and B are both well diversified. Portfolio Beta on M1 Beta on M2 Expected Return (%) A 1.5 1.9 34 B 1.8 -0.6 12 What is the expected return–beta relationship in this economy?
Answer #1

Solution:

We know that

Elrp) =ry + B1,p(E(ru) - ry] + B2,p[E(+2) - rj]

First, we calculate the risk premium, r_{p} for each of the two factors

Tp = [E(R1) -ry and

Tpz = [E(R2) -

We get the following equation to be solved

34 = 5 + 1.5 x r_{p1}+ 1.9 x r_{p2} ---------------------------------- {A}

12 = 5 + 1.8 x r_{p1} + (-0.6) x r_{p2} ------------------------------ {B}

We multiply equation A by 1.8 and B by 1.5, we get

61.2 = 9 + 2.7r_{p1} + 3.42r_{p2}

18 = 7.5 + 2.7r_{p1} -0.9 r_{p2}

- - - +

--------------------------------------------------

43.2 = 1.5 + 4.32r_{p2}

r_{p2} = 10%

Substituting the value of r_{p2} in equation A, we get

34 = 5 + 1.5 (r_{p1}) +1.9 (10)

r_{p1} = 6.67%